After a capacitor has charged for 1 tc, what percentage of current remains in the resistor?

  1. 63.2

  2. 36.8

  3. 5.0

  4. 13.5


Grade:12

2 Answers

Arun
6 years ago
Dear Abhishek
The current in the resistor has dropped to about 36.8% of the initial charging current after one time-constant of time has elapsed.
Hence option B is correct.
Regards
Arun (askIITians forum expert)
Pranjal Srivastava
6 years ago
36.8%. Since, We know that I=i(e^-t/RC) (for charging circuit)Here, R=equivalent resistance and C=euivalent capacitance I=final current i=initial current RC=time constant . For one time constant -t/RC = 1 . Therefore, I=i(e^-1) and 1/e=36.8 . Therefore, I=36.8i .therefore answer is 36.8%.
If you found this answer useful, kindly approve it.
Thankyou